# Mixed-Type Tabular Data Synthesis with Score-based Diffusion in Latent Space

### v1

@@ 4.3 Performance on Downstream Tasks
4.3.1 Machine Learning Efficiency
Real-world tabular data, such as financial data and health records, often contain some of the most sensitive and personally identifiable attributes of customers/patients. Therefore, their usage is strictly restricted (Assefa et al., 2021; Hernandez et al., 2022). Tabular data generation models address this issue by synthesizing synthetic data that conforms to the characteristics of real data but cannot be traced back to the actual data. Machine Learning Efficiency (MLE) is a crucial application for protecting data privacy. It measures whether a generative model can produce synthetic data that conforms to the characteristics of real data by comparing the performance of a machine learning model trained on real data and synthetic data. Following established settings (Kotelnikov et al., 2023; Kim et al., 2023; Lee et al., 2023), we first split a real table into a real training and a real testing set. The generative models are learned on the real training set, from which a synthetic set of equivalent size is sampled. This synthetic data is then used to train a classification/regression model (XGBoost Classifier and XGBoost Regressor (Chen & Guestrin, 2016)), which will be evaluated using the real testing set. The performance of MLE is measured by the AUC score for classification tasks and RMSE for regression tasks. The detailed settings of the MLE evaluations are in Appendix E.4.
@@ 4.3.1 Machine Learning Efficiency
Methods Adult Default Shoppers Magic Beijing News Average Gap Ranking
@@
CTGAN .886.886.886.886±.$\pm.002$696.696.696.696±.$\pm.005$875.875.875.875±.$\pm.009$855.855.855.855±.$\pm.006$902.902.902.902±.$\pm.019$880.880.880.880±.$\pm.016$ $24.5\%$ $6$
TVAE .878.878.878.878±.$\pm.004$724.724.724.724±.$\pm.005$871.871.871.871±.$\pm.006$887.887.887.887±.$\pm.003$770.770.770.770±.$\pm.011$ $1.01$±.$\pm.016$ $20.9\%$ $5$
GOGGLE .778.778.778.778±.$\pm.012$584.584.584.584±.$\pm.005$658.658.658.658±.$\pm.052$654.654.654.654±.$\pm.024$ $1.09$±.$\pm.025$877.877.877.877±.$\pm.002$ $43.6\%$ $8$
GReaT .913.913.913.913±.$\pm.003$755.755.755.755±.$\pm.006$902.902.902.902±.$\pm.005$888.888.888.888±.$\pm.008$653.653.653.653±.$\pm.013$ $13.3\%$ $4$
STaSy .906.906.906.906±.$\pm.001$752.752.752.752±.$\pm.006$914.914.914.914±.$\pm.005$934.934.934.934±.$\pm.003$656.656.656.656±.$\pm.014$871.871.871.871±.$\pm.002$ $10.9\%$ $3$
CoDi .871.871.871.871±.$\pm.006$525.525.525.525±.$\pm.006$865.865.865.865±.$\pm.006$932.932.932.932±.$\pm.003$818.818.818.818±.$\pm.021$ $1.21$±.$\pm.005$ $30.5\%$ $7$
TabDDPM .907.907.907.907±.$\pm.001$758.758.758.758±.$\pm.004$918.918.918.918±.$\pm.005$935.935.935.935±.$\pm.003$592.592.592.592±.$\pm.011$ $4.86$±$\pm 3.04$ $9.14\%^{1}$ $2$
TabSyn .$\bf{{\color[rgb]{0.8,0.25,0.33}.915}}$±.$\bf{{\color[rgb]{0.8,0.25,0.33}\pm.002}}$$\bf{{\color[rgb]{0.8,0.25,0.33}.764}}$±.$\bf{{\color[rgb]{0.8,0.25,0.33}\pm.004}}$$\bf{{\color[rgb]{0.8,0.25,0.33}.920}}$±.$\bf{{\color[rgb]{0.8,0.25,0.33}\pm.005}}$$\bf{{\color[rgb]{0.8,0.25,0.33}.938}}$±.$\bf{{\color[rgb]{0.8,0.25,0.33}\pm.002}}$$\bf{{\color[rgb]{0.8,0.25,0.33}.582}}$±.$\bf{{\color[rgb]{0.8,0.25,0.33}\pm.008}}$$\bf{{\color[rgb]{0.8,0.25,0.33}.861}}$±.$\bf{{\color[rgb]{0.8,0.25,0.33}\pm.027}}$ $\bf{{\color[rgb]{0.8,0.25,0.33}7.23\%}}$ $1$
@@
  • 2
4.3.2 Missing Value Imputation
@@ 4.4 Ablation Studies
We conduct ablation experiments to study if each component is crucial to the success of TabSyn.
@@
Figure 3: The trends of the reconstruction loss (left) and the KL-divergence loss (right) on the Adult dataset, with varying constant $\beta$, and our proposed scheduled $\beta$ ($\beta_{\rm max}=0.01,\beta_{\rm min}=10^{-5},\lambda=0.7$). The proposed scheduled $\beta$ obtains the lowest reconstruction loss with a fairly low KL-divergence loss.
Table 4: The results of single-column density and pair-wise column correlation estimation with different $\beta$ values on the Adult dataset. $\beta$ Single Pair $10^{-1}$ $1.24$ $3.05$ $10^{-2}$ $0.87$ $2.79$ $10^{-3}$ $0.72$ $2.25$ $10^{-4}$ $0.69$ $2.01$ $10^{-5}$ $41.96$ $69.17$ Scheduled $\beta$ $\mathbf{0.58}$ $\mathbf{1.54}$
Figure 3: The trends of the reconstruction loss (left) and the KL-divergence loss (right) on the Adult dataset, with varying constant $\beta$, and our proposed scheduled $\beta$ ($\beta_{\rm max}=0.01,\beta_{\rm min}=10^{-5},\lambda=0.7$). The proposed scheduled $\beta$ obtains the lowest reconstruction loss with a fairly low KL-divergence loss.
@@
Table 5: Performance of TabSyn’s variants on Adult dataset, on low-order statistics estimation tasks. Variants Single Pair TabDDPM $1.75$ $3.01$ TabSyn-OneHot $5.59$ $6.92$ TabSyn-DDPM $1.02$ $2.15$ TabSyn $\mathbf{0.58}$ $\mathbf{1.54}$
@@ 5 Conclusions and Future Directions
For future research, one potential direction is exploring the application of TabSyn in other scenarios, such as conditional generation or controllable generation. In addition, TabSyn adopts an MLP model as the denoising function, which is simple and already demonstrates competitive performance. It would be interesting to study if more advanced designs of the denoising function can lead to better generation quality.
@@ E.4 Details of Machine Learning Efficiency Experiments
As preliminarily illustrated in Sec. 4.3.1 and Appendix E.1, each dataset is first split into the real training and testing set. The generative models are learned on the real training set. After the models are learned, a synthetic set of equivalent size is sampled.
@@ F.2 Sample-wise Quality score of Synthetic Data ($\alpha$-Precison and $\beta$-Recall)
F.3 Missing Value Imputation

### v3

@@ 4.3 Performance on Downstream Tasks
Machine Learning Efficiency.
We then evaluate the quality of synthetic data by evaluating their performance in Machine Learning Efficiency tasks. Following established settings (Kotelnikov et al., 2023; Kim et al., 2023; Lee et al., 2023), we first split a real table into a real training and a real testing set. The generative models are learned on the real training set, from which a synthetic set of equivalent size is sampled. This synthetic data is then used to train a classification/regression model (XGBoost Classifier and XGBoost Regressor (Chen & Guestrin, 2016)), which will be evaluated using the real testing set. The performance of MLE is measured by the AUC score for classification tasks and RMSE for regression tasks. The detailed settings of the MLE evaluations are in Appendix E.4.
@@
Methods Adult Default Shoppers Magic Beijing News1 Average Gap
@@
SMOTE .899.899.899.899±.$\pm.007$741.741.741.741±.$\pm.009$911.911.911.911±.$\pm.012$934.934.934.934±.$\pm.008$593.593.593.593±.$\pm.011$897.897.897.897±.$\pm.036$ $9.39\%$
CTGAN .886.886.886.886±.$\pm.002$696.696.696.696±.$\pm.005$875.875.875.875±.$\pm.009$855.855.855.855±.$\pm.006$902.902.902.902±.$\pm.019$880.880.880.880±.$\pm.016$ $24.5\%$
TVAE .878.878.878.878±.$\pm.004$724.724.724.724±.$\pm.005$871.871.871.871±.$\pm.006$887.887.887.887±.$\pm.003$770.770.770.770±.$\pm.011$ $1.01$±.$\pm.016$ $20.9\%$
GOGGLE .778.778.778.778±.$\pm.012$584.584.584.584±.$\pm.005$658.658.658.658±.$\pm.052$654.654.654.654±.$\pm.024$ $1.09$±.$\pm.025$877.877.877.877±.$\pm.002$ $43.6\%$
GReaT .913.913.913.913±.$\pm.003$755.755.755.755±.$\pm.006$902.902.902.902±.$\pm.005$888.888.888.888±.$\pm.008$653.653.653.653±.$\pm.013$ OOM $13.3\%$
STaSy .906.906.906.906±.$\pm.001$752.752.752.752±.$\pm.006$914.914.914.914±.$\pm.005$934.934.934.934±.$\pm.003$656.656.656.656±.$\pm.014$871.871.871.871±.$\pm.002$ $10.9\%$
CoDi .871.871.871.871±.$\pm.006$525.525.525.525±.$\pm.006$865.865.865.865±.$\pm.006$932.932.932.932±.$\pm.003$818.818.818.818±.$\pm.021$ $1.21$±.$\pm.005$ $30.5\%$
TabDDPM2 .907.907.907.907±.$\pm.001$758.758.758.758±.$\pm.004$918.918.918.918±.$\pm.005$935.935.935.935±.$\pm.003$592.592.592.592±.$\pm.011$ $4.86$±$\pm 3.04$ $9.14\%^{1}$
TabSyn .$\bf{{\color[rgb]{0.8,0.25,0.33}.915}}$±.$\bf{{\color[rgb]{0.8,0.25,0.33}\pm.002}}$$\bf{{\color[rgb]{0.8,0.25,0.33}.764}}$±.$\bf{{\color[rgb]{0.8,0.25,0.33}\pm.004}}$$\bf{{\color[rgb]{0.8,0.25,0.33}.920}}$±.$\bf{{\color[rgb]{0.8,0.25,0.33}\pm.005}}$$\bf{{\color[rgb]{0.8,0.25,0.33}.938}}$±.$\bf{{\color[rgb]{0.8,0.25,0.33}\pm.002}}$$\bf{{\color[rgb]{0.8,0.25,0.33}.582}}$±.$\bf{{\color[rgb]{0.8,0.25,0.33}\pm.008}}$$\bf{{\color[rgb]{0.8,0.25,0.33}.861}}$±.$\bf{{\color[rgb]{0.8,0.25,0.33}\pm.027}}$ $\bf{{\color[rgb]{0.8,0.25,0.33}7.23\%}}$
@@
Missing Value Imputation.
@@ 4.4 Ablation Studies
Figure 3: The trends of the validation reconstruction (left) and KL-divergence (right) losses on the Adult dataset, with varying constant $\beta$, and our proposed scheduled $\beta$ ($\beta_{\rm max}=0.01,\beta_{\rm min}=10^{-5},\lambda=0.7$). The proposed scheduled $\beta$ obtains the lowest reconstruction loss with a fairly low KL-divergence loss.
Table 4: The results of single-column density and pair-wise column correlation estimation with different $\beta$ values on the Adult dataset.
$\beta$ Single Pair
$10^{-1}$ $1.24$ $3.05$
$10^{-2}$ $0.87$ $2.79$
$10^{-3}$ $0.72$ $2.25$
$10^{-4}$ $0.69$ $2.01$
$10^{-5}$ $41.96$ $69.17$
Scheduled $\beta$ $\mathbf{0.58}$ $\mathbf{1.54}$
@@
Table 5: Performance of TabSyn’s variants on Adult dataset, on low-order statistics estimation tasks.
Variants Single Pair
TabDDPM $1.75$ $3.01$
TabSyn-OneHot $5.59$ $6.92$
TabSyn-DDPM $1.02$ $2.15$
TabSyn $\mathbf{0.58}$ $\mathbf{1.54}$
@@ E.4 Details of Machine Learning Efficiency Experiments
As preliminarily illustrated in Sec. 4.3 and Appendix E.1, each dataset is first split into the real training and testing set. The generative models are learned on the real training set. After the models are learned, a synthetic set of equivalent size is sampled.
@@ F.2 Sample-wise Quality score of Synthetic Data ($\alpha$-Precison and $\beta$-Recall)
F.3 Detection: Classifier Two Sample Tests (C2ST)
We further study how difficult it is to tell apart the real data from the synthetic data, therefore evaluating if the synthetic data can recover the real data distribution. To this end, we employ the detection metric provided by sdmetrics  [^9] https://docs.sdv.dev/sdmetrics/metrics/metrics-in-beta/detection-single-table. In Table 11, we present the results obtained using logistic regression as the detection method.
Table 11: Detection score (C2ST) using logistic regression classifier. Higher scores indicate better performance.
Method Adult Default Shoppers Magic Beijing News
SMOTE $0.9710$ $0.9274$ $0.9086$ $0.9961$ $0.9888$ $0.9344$
CTGAN $0.5949$ $0.4875$ $0.7488$ $0.6728$ $0.7531$ $0.6947$
TVAE $0.6315$ $0.6547$ $0.2962$ $0.7706$ $0.8659$ $0.4076$
GOGGLE $0.1114$ $0.5163$ $0.1418$ $0.9526$ $0.4779$ $0.0745$
GReaT $0.5376$ $0.4710$ $0.4285$ $0.4326$ $0.6893$
STaSy $0.4054$ $0.6814$ $0.5482$ $0.6939$ $0.7922$ $0.5287$
CoDi $0.2077$ $0.4595$ $0.2784$ $0.7206$ $0.7177$ $0.0201$
TabDDPM $0.9755$ $0.9712$ $0.8349$ $\bf{{\color[rgb]{0.8,0.25,0.33}0.9998}}$ $0.9513$ $0.0002$
TabSyn $\bf{{\color[rgb]{0.8,0.25,0.33}0.9986}}$ $\bf{{\color[rgb]{0.8,0.25,0.33}0.9870}}$ $\bf{{\color[rgb]{0.8,0.25,0.33}0.9740}}$ $0.9732$ $\bf{{\color[rgb]{0.8,0.25,0.33}0.9603}}$ $\bf{{\color[rgb]{0.8,0.25,0.33}0.9749}}$
As indicated in the table, the Detection score exhibits superior discriminative power compared to other metrics, such as single-column density estimation, pair-wise column shape estimation, and MLE. The detection score shows significant variations across different models for synthetic data generation. As indicated in the table, the Detection score exhibits superior discriminative power compared to other metrics, such as single-column density estimation, pair-wise column shape estimation, and MLE. The detection score shows significant variations across different models for synthetic data generation. The proposed TabSyn consistently achieves notably high scores across all datasets (SMOTE (Chawla et al., 2002) directly interpolates within the training set, so it is not surprising that it achieves high scores in the detection metric.).
F.4 Missing Value Imputation
@@
F.5 Impacts of the quality of VAEs
Intuitively, the performance of TabSyn appears to be highly dependent on the quality of the pre-trained VAE model. Therefore, we conduct further research to investigate how a less trained VAE model would impact the quality of synthetic data generated by TabSyn. To this end, we investigate the quality of synthetic data generated by TabSyn using the embeddings of the VAE obtained at different epochs as the latent space. Figure 9 plots the results of single-column density estimation and pair-wise column correlation estimation on the Adult and Default datasets, with intervals set at 400 epochs. We can observe that increasing the training epochs of the VAE does indeed improve the quality of TabSyn’s generated data. Additionally, even when the VAE is sub-optimal (e.g., training epochs around 2000), TabSyn’s performance is already very close to the optimal ones. Furthermore, even with a relatively low number of VAE training epochs (e.g., 800-1200), TabSyn’s performance approaches or even surpasses the most competitive baseline, TabDDPM. Based on this observation, we recommend thoroughly training the VAE to achieve superior data generation quality when resources are abundant. However, when resources are limited, reducing the VAE training duration still yields decent performance.
Figure 9: Performance of TabSyn with VAEs trained for different epochs. By default, TabSyn’s VAE is trained with 4000 epochs
F.6 Privacy Protection: Distance to Closest Record (DCR)
To study if the synthetic data will cause privacy information leakage issues, we compute the DCRs of the synthetic data. Specifically, we follow the ’synthetic vs. holdout’ setting  [^10] https://www.clearbox.ai/blog/2022-06-07-synthetic-data-for-privacy-
preservation-part-2
. We initially divide the dataset into two equal parts: the first part served as the training set for training our generative model, while the second part was designated as the holdout set, which is not used for training. After completing model training, we sample a synthetic set of the same size as the training set (and the holdout set).
We then calculate the DCR scores for each sample in the synthetic set concerning both the training set and the holdout set. We can create histograms to visualize the distribution of DCR scores for the synthetic set in comparison to both the training and holdout sets. Intuitively, if there is a privacy issue (e.g. if the synthetic set is directly copied from the training set), then the DCR scores for the training set should be closer to 0 than those for the testing set. Conversely, if there is no privacy issue, the distributions of DCR scores of the training and holdout sets should largely overlap. In Figure 10, we plot the distributions of synthetic sets’ DCRs concerning the training set and holdout set on Default and Shoppers. We can observe that deep generative models such as CoDi, STaSy, TabDDPM, and TabSyn do not suffer from privacy issues, while the interpolation-based method SMOTE might not be able to protect privacy information.
Figure 10: Distributions of the DCR scores between the synthetic dataset and the training/holdout datasets. Deep generative models have similar DCR distributions concerning the training set and holdout set, while in the interpolation-based method SMOTE, DCRs concerning the training set are much smaller than DCRs concerning the holdout set.
Additionally, we calculate the probability that a synthetic sample is closer to the training set (rather than the holdout set). When this probability is close to 50% (i.e., 0.5), it indicates that the distribution of distances between synthetic and training instances is very similar (or at least not systematically smaller) than the distribution of distances between synthetic and holdout instances, which is a positive indicator in terms of privacy risk. Table 13 displays the results obtained by different models on Default and Shoppers datasets.
Table 13: DCR score (the probability that a synthetic example is closer to the training set rather than the holdout set (%, a score closer to $50\%$ is better).
Method Default Shoppers
SMOTE $91.41$%±$\pm 3.42$ $96.40$%±$\pm 4.70$
STaSy $50.23$%±$\pm 0.09$ $51.53$%±$\pm 0.16$
CoDi $51.82$%±$\pm 0.26$ $51.06$%±$\pm 0.18$
TabDDPM $52.15$%±$\pm 0.20$ $63.23$%±$\pm 0.25$
TabSyn $51.20$%±$\pm 0.18$ $52.90$% ±$\pm 0.22$